# Mesh Watermark Removal Attack and Mitigation: A Novel Perspective of Function Space

### v2

@@ 4.1 Removal Implementation
$\mathcal{L}_{F_{\Theta}}$ $=\int_{\mathbf{x}\in V_{0}\cup V_{1}}\lambda_{1}\lVert F_{\Theta}(\mathbf{x})-F(\mathbf{x})\rVert+\lambda_{2}\lVert|\nabla F_{\Theta}(\mathbf{x})|-1\rVert d\mathbf{x}$ (4)
@@
where $\lambda_{1}=3000,\lambda_{2}=5,\lambda_{3}=100$. For on-surface points $V_{0}$, their ground truth signed distance values $F(\mathbf{x})$ and first-order derivatives $\nabla F_{\Theta}(\mathbf{x})$ are assigned with zero and their normal vectors, respectively. For off-surface points $V_{1}$, their ground truth signed distance values are assigned with values obtained in the sampling process. We use SIREN (Sitzmann et al. 2020) as the backbone of $F_{\Theta}$, which consists of four Linear layers with their channel size set to 512. We use Adam optimizer with the initialized learning rate $10^{-3}$ for 1000 epochs, and we decrease the learning rate by half every 200 epochs. After obtaining the fitted SDF, the final step is to obtain the post-processed mesh $\mathcal{M}^{\prime}_{w}$ by extracting it from SDF through isosurfacing. Unless mentioned explicitly, the default isosurfacing method is marching cube (Lorensen and Cline 1998).
@@ 4.2 Removal Evaluation
Dataset, Metrics and Examined Removal Methods
Metrics and Datasets. We evaluate the evasion rate for the effectiveness of watermark removal and Hausdorff distance for mesh similarity between the watermarked mesh $\mathcal{M}_{w}$ and the post-processed mesh $\mathcal{M}^{\prime}_{w}$. Let $w^{\prime}$ be the message extracted from $\mathcal{M}^{\prime}_{w}$. We judge the removal method successfully evades the detection if $BA(w,w_{A})<\frac{\tau}{n}$. We select the message length $n=48$ and $\tau=31$ in our experiment. In this case, we have a false positive rate $FPR(\tau)<0.05$. We normalize meshes in ShapeNet (Chang et al. 2015) and Stanford Repo (Laboratory 2023) to a unit sphere. All watermark embedding, extraction, and removal experiments are conducted on normalized meshes.
@@
Examined post-processed methods. We first examine popular mesh processing methods such as Gauss noise, rotation, quantization, and smoothing, which are commonly used in previous works (Wang et al. 2022; Zhu et al. 2024) to evaluate the robustness of their watermark. These post-processed methods output $\mathcal{M}^{\prime}_{w}$ sharing the same topology with $\mathcal{M}_{w}$. Our FuncEvade and remeshing (Botsch and Kobbelt 2004; Botsch et al. 2010) will output a $\mathcal{M}^{\prime}_{w}$ with different topology with $\mathcal{M}_{w}$. In detail, remeshing incrementally performs simple operations such as edge splits, edge collapses, edge flips, and Laplacian smoothing. All the vertices of the remeshed patch are reprojected to the original surface to keep an approximation of the input. We set the target edge length of remeshing to be half of the average edge length of the input mesh. For other removal methods, we use Gauss noise with mean $\mu=0$ and standard deviation $\sigma=0.005$, rotation with rotation axis $(1,0,0)$ and rotation angle $\alpha<\frac{\pi}{3}$, quantization with bits $N_{b}=4$, laplacian smoothing with $\lambda=2$ and smoothing iteration $10$.
@@ 5.2 Watermark Embedding
Newton’s method calculates $D^{-1}(\mathbf{x})$. We derive $\mathbf{y}=D^{-1}(\mathbf{x})$ by calcuating the zero point of $\mathbf{y}\pm\nabla F_{\Theta}(\mathbf{y})C(y)-\mathbf{x}=0$. We apply Newton’s method (Ypma 1995) to calculate the zero point as the following repeated process:
@@
where $\mathbf{y}_{i}$ is the guessed zero point at $i$-th iteration, $\cdot$ is the matrix multiplication operation, and $\mathcal{J}^{-1}_{D}(\mathbf{y})$ is the inverse of Jacobian matrix of $D(\mathbf{y})$. To calculate a correct zero point, this process is sensitive to the accuracy of $\mathcal{J}_{D}(\mathbf{y})$, which is derived from Hessian of the SDF: $\mathcal{H}_{F_{\Theta}}(\mathbf{y})$. However, $\nabla F_{\Theta}(\mathbf{y})$ is supervised with ground truth value only when $\mathbf{y}$ is on zero-isosurface, which means $\nabla F_{\Theta}(\mathbf{y})$ may not be accurate when $\mathbf{y}$ is far from surface. Thus, Newton’s method may not find a zero point if $y_{0}$ is initialized far from the surface. To address this issue, we randomly initialize multiple $y_{0}$ within $[-1,1]^{3}$. One hundred samples are enough to find the ground truth zero point.
@@ 5.4 Watermark Evaluation
Setup
Since we embed each spherical partition with one binary bit in the secret message $w$, if the number of partitions $N_{s}*N_{s}$ is greater than the length of the message, we repeatedly embed message bits until we consume all partitions. Unless explicitly mentioned, our experiment is conducted under the following settings. We set $N_{s}=32$ (i.e., the spherical system is divided into $32*32$ partitions). We set message length $n=48$, and the detection threshold $\tau=31$. We use marching cube (Lorensen and Cline 1998) as our default isosurfacing method and the default watermarking strength $\delta=0.001$. We use the same dataset, evaluation metrics, and watermarking methods (as baselines) in Sec. 4.2 to evaluate watermark accuracy before any post-processing method is applied. We use the same set of attack settings for all post-processing methods in Sec. 4.2 to compare evasion rates between baselines and FuncMark.

### v3

@@ 4.1 Removal Implementation
$\int_{\mathbf{x}\in V_{0}\cup V_{1}}\lambda_{2}\lVert F_{\Theta}(\mathbf{x})-F(\mathbf{x})\rVert+\lambda_{3}\lVert|\nabla F_{\Theta}(\mathbf{x})|-1\rVert d\mathbf{x}$
where $\lambda_{1}=100,\lambda_{2}=3000,\lambda_{3}=5$. For on-surface points $V_{0}$, their ground truth signed distance values $F(\mathbf{x})$ and first-order derivatives $\nabla F_{\Theta}(\mathbf{x})$ are assigned with zero and their normal vectors, respectively. For off-surface points $V_{1}$, their ground truth signed distance values are assigned with values obtained in the sampling process. We use SIREN (Sitzmann et al. 2020) as the backbone of $F_{\Theta}$, which consists of four Linear layers with their channel size set to 512. We use Adam optimizer with the initialized learning rate $10^{-3}$ for 1000 epochs, and we decrease the learning rate by half every 200 epochs. After obtaining the fitted SDF, the final step is to obtain the post-processed mesh $\mathcal{M}^{\prime}_{w}$ by extracting it from SDF through marching cube (Lorensen and Cline 1998).
@@ 4.2 Removal Evaluation
Metrics and Datasets. We evaluate the evasion rate for the effectiveness of watermark removal and Hausdorff distance for mesh similarity between the watermarked mesh $\mathcal{M}_{w}$ and the post-processed mesh $\mathcal{M}^{\prime}_{w}$. Let $w^{\prime}$ be the message extracted from $\mathcal{M}^{\prime}_{w}$. We judge the removal method successfully evades the detection if $BA(w,w_{A})<\frac{\tau}{n}$. We select the message length $n=48$ and $\tau=31$ in our experiment. In this case, we have a false positive rate $FPR(\tau)<0.05$. We normalize meshes in ShapeNet (Chang et al. 2015) and Stanford Repo (Laboratory 2023) to $[-1,1]^{3}$. All watermark embedding, extraction, and removal experiments are conducted on normalized meshes.
@@
Examined post-processed methods. Following the above watermark process, we post-process watermarked mesh to remove the watermark. We first examine popular mesh processing methods such as Gauss noise, rotation, quantization, and smoothing, which are commonly used in previous works (Wang et al. 2022; Zhu et al. 2024) to evaluate the robustness of their watermark. These post-processed methods output $\mathcal{M}^{\prime}_{w}$ sharing the same topology with $\mathcal{M}_{w}$. Our FuncEvade and remeshing (Botsch and Kobbelt 2004; Botsch et al. 2010) will output a $\mathcal{M}^{\prime}_{w}$ with different topology with $\mathcal{M}_{w}$. In detail, remeshing incrementally performs simple operations such as edge splits, edge collapses, edge flips, and Laplacian smoothing. All the vertices of the remeshed patch are reprojected to the original surface to keep an approximation of the input. We set the target edge length of remeshing to be half of the average edge length of the input mesh. For other removal methods, we use Gauss noise with mean $\mu=0$ and standard deviation $\sigma=0.005$, rotation with rotation axis $(1,0,0)$ and rotation angle $\alpha<\frac{\pi}{3}$, quantization with bits $N_{b}=6$, laplacian smoothing with $\lambda=2$ and smoothing iteration $10$.
@@ 5.2 Watermark Embedding
Newton’s method calculates $D^{-1}(\mathbf{x})$. We derive $\mathbf{y}=D^{-1}(\mathbf{x})$ by calcuating the zero point of $\mathbf{y}\pm\delta\nabla F_{\Theta}(\mathbf{y})-\mathbf{x}=0$. We apply Newton’s method (Ypma 1995) to calculate the zero point as the following repeated process:
@@
where $\mathbf{y}_{i}$ is the guessed zero point at $i$-th iteration, $\cdot$ is the matrix multiplication operation, and $\mathcal{J}^{-1}_{D}(\mathbf{y})$ is the inverse of Jacobian matrix of $D(\mathbf{y})$. To calculate a correct zero point, this process is sensitive to the accuracy of $\mathcal{J}_{D}(\mathbf{y})$, which is derived from Hessian of the SDF: $\mathcal{H}_{F_{\Theta}}(\mathbf{y})$. However, $\nabla F_{\Theta}(\mathbf{y})$ is supervised with ground truth value only when $\mathbf{y}$ is on zero-isosurface, which means $\nabla F_{\Theta}(\mathbf{y})$ may not be accurate when $\mathbf{y}$ is far from surface. Thus, Newton’s method may not find a zero point if $\mathbf{y}_{0}$ is initialized far from the surface. To address this issue, we randomly initialize multiple $\mathbf{y}_{0}$ within $[-1,1]^{3}$. One hundred samples are enough to find the ground truth zero point.
@@ 5.4 Watermark Evaluation
Setup. Since we embed each spherical partition with one binary bit in the secret message $w$, if the number of partitions $N_{s}*N_{s}$ is greater than the length of the message, we repeatedly embed message bits until we consume all partitions. Unless explicitly mentioned, our experiment is conducted under the following settings. We set $N_{s}=32$ (i.e., the spherical system is divided into $32*32$ partitions). We set message length $n=48$, and the detection threshold $\tau=31$. We use marching cube (Lorensen and Cline 1998) as our default isosurfacing method and the default watermarking strength $\delta=0.001$. We use the same dataset, evaluation metrics, and watermarking methods (as baselines) in Sec. 4.2 to evaluate watermark accuracy before any post-processing method is applied. We use the same set of attack settings for all post-processing methods in Sec. 4.2 to compare evasion rates between baselines and FuncMark.